OpenStudy (anonymous):

√99 simplify each radical expression

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{99}\] \[\Large \sqrt{9*11}\] \[\Large \sqrt{9}*\sqrt{11}\] \[\Large 3\sqrt{11}\] So \[\Large \sqrt{99} = 3\sqrt{11}\]

OpenStudy (anonymous):

√225

OpenStudy (anonymous):

√225 √25 * 9 5 * 3 15

OpenStudy (anonymous):

√128

OpenStudy (anonymous):

You're not going to learn this way.

OpenStudy (anonymous):

8√2

OpenStudy (anonymous):

I go to tutoring but the teacher won't help

OpenStudy (anonymous):

The first answer showed you how to approach this problem: try to express the number under the square root as multiplication of two numbers, at least one of which is a square. Then separate and simplify, using: \[\sqrt{x*y} = \sqrt{x} * \sqrt{y}\]

OpenStudy (anonymous):

With \[\sqrt{128}\] you should have asked yourself "what two numbers multiplied by each other give 128"?

OpenStudy (anonymous):

For even numbers, it's easy to just continually divide by 2. If you can not find another number off the top of your head. But Knowak is right with what he posted.